Grade 12th passIIT JEE Entrance Exam

a particle starts with velocity 10m/s the acceleration of the particle is 2m/s^2 find
1.the velocity of the particle at t=4?
2.find displacement of the particle in 2s?
3.the displacement of the particle in 2nd second is?

Shaila
8 Years agoGrade 12th pass

1 Answer

Eshan
ApprovedApproved Tutor Answer8 Years ago
Dear student,

Using the equations of motion,

v=u+at=10+(2)(4)=18m/s
Displacment in 2 seconds=s_2=ut+\dfrac{1}{2}at^2=(10)(2)+\dfrac{1}{2}(2)(2^2)=24m
Displacement in 1 seocond=s_1=(10)(1)+\dfrac{1}{2})(2)(1^2)=11m

Displacement in 2ndsecond= Displacment in 2 seconds-Displacement in 1 second
=s_2-s_1=13m